if one of the solution of the equation x^3-2x^2+ax+10=0 is the additive inverse of another,then which inequality is true? a) -40b) -10

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
 let one solution be m, then the other solution will be -m.
       let third solution be n
       sum of roots taken one at a time= m+(-m)+n= n = 2 -------- (A)
       sum of roots taken two at a time = -m2 -mn+mn = -m2 = a ------(B)
       product of roots = -m2n = -10------(C)
 from A, B, C we get:
   a= -5 .
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)